A 2.59 g sample of a compound consisting of carbon, hydrogen, oxygen, nitrogen, and sulfur was combusted in excess oxygen. This produced 3.63 g CO2 and 0.444 g H2O . A second sample of this compound with a mass of 4.89 g produced 2.49 g SO3 . A third sample of this compound with a mass of 7.36 g produced 2.95 g HNO3 . Determine the empirical formula of the compound. Enter the correct subscripts on the given chemical formula.
1
Expert's answer
2019-11-05T07:40:30-0500

n(CO2) = 3.63g/44g/mol = 0.0825 mol

n(SO3) = 0.031125 mol

n(HNO3) = 0.046825 mol

M = m/n = 4.89g/0.031125mol = 157 g/mol

Empirical formula of the compound C5H20O2SN